The two ends of a train moving with constant acceleration pass a certain point with velocity u and v. The velocity with which the middle point of the train passes the same point is?

Grade:12th pass

2 Answers

Bejoi Mathew
7 years ago
We all know that,
v2-u2=2as or v2=u2+2as -----> (1)
a=v2-u2/2s
Now for the mid-point of the train
s1=s/2
a1=v12-u2/2s1
a1=v12-u2/2s/2
a1=v12-u2/s or v12=u+ as ----> (2)
from (1) and (2) you will get
as= v2-u2/2
v12=u+ as
v12= v2+u/2
HERE IS YOUR ANSWER HOPE THIS HELPS!
THANKS 
 
 
 
 
 
ankit singh
3 years ago
Two ends of a train moving with a constant acceleration pass a certain point with velcities uandv. Show that the velocity with which the middle point of the train passes the same point is `sqrt(9u"^(2) +v^(2)//2). or V=√(u2+v2)/2.